# Shadow-FT: Tuning Instruct via Base

### v2

@@ Abstract
Large language models (LLMs) consistently benefit from further fine-tuning on various tasks. However, we observe that directly tuning the Instruct (i.e., instruction tuned) models often leads to marginal improvements and even performance degeneration. Notably, paired Base models, the foundation for these Instruct variants, contain highly similar weight values (i.e., less than 2% on average for Llama 3.1 8B). Therefore, we propose a novel Shadow-FT framework to tune the Instruct models by leveraging corresponding Base models. The key insight is to fine-tune the Base model, and then directly graft the learned weight updates to the Instruct model. Our proposed Shadow-FT introduces no additional parameters, is easy to implement, and significantly improves performance. We conduct extensive experiments on tuning mainstream LLMs, such as Qwen 3 and Llama 3 series, and evaluate them across 19 benchmarks covering coding, reasoning, and mathematical tasks. Experimental results demonstrate that Shadow-FT consistently outperforms conventional full-parameter and parameter-efficient tuning approaches. Further analyses indicate that Shadow-FT can be applied to multimodal large language models (MLLMs) and combined with direct preference optimization (DPO). Codes and weights are available at Github.
@@ 1 Introduction
Large Language Models (LLMs), such as Qwen [2], Llama [1], and Gemma [3], have demonstrated remarkable performance across diverse disciplines [4, 7]. Such a strong capability is always attributed to the pre-training on massive data with billions of parameters [9, 10]. When applied in real-world scenarios, there are several challenges. The users want the LLMs to follow their instructions helpfully and honestly [11], which is not covered during the pre-training [4, 8]. Meanwhile, the downstream tasks always involve specific domain knowledge requiring adaptation [12, 13].
@@
To tackle these issues, one predominant approach is further tuning LLMs on desired tasks, including full parameter fine-tuning and parameter-efficient fine-tuning [15, 14]. Typically, for each model size, two paired variants are provided: the pretrained base model (denoted as Base) and its instruction-tuned version (denoted as Instruct). The Base model exhibits relatively poor instruction-following ability, while the Instruct model performs better. However, we observe that tuning the Instruct models brings marginal improvements and even a performance degeneration. Therefore, how to tune the Instruct model effectively gains increasing importance.
@@
Figure 1: The similarity between Base and Instruct models (part a), vanilla SFT on Instruct (part b), and the Shadow-FT framework (part c). Based on the observations that the two models are highly similar (gap $\sigma$ less than 0.02), we propose to tune the paired Base model and then graft the weight updates onto Instruct model.
@@
In this paper, we first analyze the weights of paired Base and Instruct models considering the relative absolute difference $\sigma$. Fortunately, we find that the weights of Base and Instruct are highly similar. As shown in Figure 1, the gap $\sigma$ is quite low, such as an average $\sigma$ of 0.016 for the Llama-3.1-8B model. Intuitively, the contained instruction following ability of Instruct model might disturb the learning of new knowledge, while Base can avoid it. Motivated by these, we thus propose a novel Shadow-FT framework to employ the Base model as ’shadow’ of Instruct. The key is to tune the Base for better weight updates and directly graft these updates to Instruct, as they share the same structures.
@@
To evaluate the performance, we conduct extensive experiments tuning mainstream LLMs such as Qwen 3 [2] and Llama 3 [1]. For the tuning data, we employ the BAAI-Infinity-Instruct Dataset [^1] https://huggingface.co/datasets/BAAI/Infinity-Instruct and extract 2000 samples named as BAAI-2k following [17, 18]. Without the loss of generality, we apply Shadow-FT on full parameter and low-rank settings, and then report the performance on 19 datasets. Experimental results indicate that Shadow-FT consistently outperforms the baselines under various settings, demonstrating the effectiveness and robustness. Further analyses show that Shadow-FT can be applied to MLLMs and combined with DPO for alignment. Our contributions can be concluded as follows:
@@ 2.1 Background
Supervised Fine-tuning (SFT) is a fundamental approach to updating the knowledge of LLMs. Vanilla SFT methods update all the parameters by gradient descent following $W^{+}\leftarrow W+\Delta W$, where $W\in\mathbb{R}^{d_{1}\times d_{2}}$ is an arbitrary weight and $W^{+}$ is the updated variant. To reduce the update costs, LoRA [14] introduces a low-rank branch to learn the weight updates following $W^{+}\leftarrow W+AB$, where $A\in\mathbb{R}^{d_{1}\times r},B\in\mathbb{R}^{r\times d_{2}}$ and $r\ll\min\{d_{1},d_{2}\}$. The original weight $W$ is frozen during training, and only the low-rank branch is updated.
@@
Current LLMs typically follow a two-stage training pipeline, including pre-training and post-training. During pre-training, LLMs are trained on massive training data on next token prediction tasks [31], and the weights would be released as Base version. The Instruct variant, post-trained upon the Base model, is further aligned with human preference and tuned for reasoning tasks [20]. Therefore, Instruct model performs better than Base model regarding instruction-following ability.
@@ 2.2 Directly Tuning Instruct
However, tuning the Instruct models often leads to marginal improvements and even performance degeneration. Table 1 shows the tuned performance of the Instruct models using the BAAI-2k. We report the average scores of popular benchmarks. Compared to the vanilla Instruct, the tuned version shows marginal improvement, and even degeneration in more cases. Specifically, as shown in Table 1, tuning Qwen-3-4B on BAAI-2k dataset via conventional LoRA would lead to a drop of 2.6 in Math-7 (from 73.8 to 71.2), 6.8 in Code-3 (from 66.4 to 59.6), and 2.6 in Reasoning-9 (from 63.7 to 61.1). Therefore, how to effectively tune Instruct remains a challenge.
@@ 2.3 Similar Weights: Base & Instruct
where $\sum$ is the element-wise sum and ||$\cdot$|| ⋅ | means the absolute operations. The $\sigma$ would be 1 if one is much larger than the other, and be 0 if the two matrices are exactly the same. The smaller the $\sigma$, the more similar the two matrices are. Figure 2 shows the weights of the same layer from Base and Instruct, and also their differences with $\sigma=0.016$. We can find that the gaps are quite small and negligible after zooming in 10x in the z-axis. Please refer to Appendix A for more $\sigma$ regarding various LLMs. In summary, these paired Base and Instruct models are highly similar with $\sigma<0.02$.
@@ 3.1 Shadow-FT
where $Tune$ is the fine-tuning method, such as full-parameter fine-tuning and LoRA. After that, we would like to get the weights updates as the learned knowledge using, and directly graft these updates to the Instruct model as:
@@ 3.2 Relation with Task Vectors
Task Vectors aim to represent the ability on tasks as vectors, and are widely used for arithmetic operations on these tasks regarding the same base model [34]. Chat Vector [33] extends such an idea to LLMs, which models weight differences between Instruct and Base models as vectors and then adds the vectors to continually pretrained Base models. Specifically, Chat Vector continually pre-trains Llama2 [35] on the Traditional Chinese corpus, and then adds on the chat vectors. Compared to Chat Vector [33], the differences are as follows: 1) task: Chat Vector focuses on continual pertaining while Shadow-FT can be applied to board tuning methods, including full-parameter fine-tuning, LoRA, and DPO. 2) motivation: Chat Vector aims to extend the language ability. Shadow-FT aims to tackle the degeneration issue based on the weight similarity.
@@ 4.1 Experimental Setup
For the tuning data, we build BAAI-2k by extracting 2000 samples from BAAI-Infinity-Instruct Dataset  [^2] https://huggingface.co/datasets/BAAI/Infinity-Instruct/tree/main/Gen following [17, 18]. We select the samples with high rewards to ensure the data quality and uniform sampling among all categories for data diversity. Without loss of generality, we tune various LLMs, including Qwen 3 series [39] and Llama 3 series [1]. Also, we report the results on Gemma-3 series [3], Yi series [69], and Falcon series [70] in Section 5.5. We employ LLaMA-Factory [36] for the code base and apply two tuning strategies: full-parameter and LoRA. All experiments are conducted on 8 A100 GPUs. Please refer to Appendix B.1 for detailed hyperparameters.
@@
To evaluate the tuned LLMs on downstream benchmarks, we employ the OpenCompass framework [37] and lmdeploy as the acceleration framework [38]. During inference, we set the cutoff length as 4096 and the batch size as 512. Considering the benchmarks, we select three representative abilities, i.e., mathematical, coding, and commonsense reasoning ability, and report the average scores marked as Math-7, Code-3, and Reasoning-9. Specifically, Math-7 denotes the results of AIME24 [5], GSM8K (0-shot and 8-shot) [6], MATH [63], MATH-500, Minerva_Math [65], SVAMP [59]. Code-3 for HumanEval [52], HumanEval+ [53], LiveCodeBench [58]. Reasoning-9 for ARC-challenge [55], BBH (0-shot and few-shot), DROP [51], GPQA Diamond [60], MMLU [56], MMLU Pro [47], Winogrande [57], TheoremQA [54]. To avoid the impact of different prompts, we mainly evaluate under a zero-shot setting. Please refer to Appendix B.2 for more details.
@@ 4.2 Main Results
Table 1: Performance comparison of different methods tuning popular LLMs. Math-7 denotes the average score of 7 mathematical benchmarks including AIME24, Code-3 for 3 code benchmarks including LiveCodeBench, and Reasoning-9 for 9 commonsense reasoning benchmarks including MMLU Pro. For Math-7 and Code-3, we report the mean value of three runs. We employ the Instruct version and report the final average scores. Please refer to Appendix B.3 for detailed scores.
Model Method Math-7 Code-3 Reasoning-9 Avg.
@@
  • While conventional tuning fails, Shadow-FT performs well in adaptation at the same cost. Across all model sizes and tasks, Shadow-FT consistently outperforms tuning baselines and vanilla Instruct model. For example, on Qwen-3-4B, Shadow-FT archives an average score of 69.4, which is 3.7 higher than the 65.7 of conventional tuning methods and 1.4 higher than the vanilla Instruct model. The conclusion is consistent on larger models such as Qwen-3-14B. Moreover, Shadow-FT does not introduce any extra training overheads. These consistent gains demonstrate that our proposed Shadow-FT can effectively learn the knowledge contained in training data.
  • Shadow-FT works well under both full-parameter setting and LoRA. For instance, when tuning Qwen-3-4B under full-parameter setting, Shadow-FT achieves 73.7/67.4/64.9 on Math-7/Code-3/Reasoning-9 compared to 72.9/66.4/62.9 of conventional tuning methods. When applying a low-rank setting, Shadow-FT achieves 75.9/69.7/65.0, which is 4.7/10.1/3.9 higher than conventional LoRA. These indicate that Shadow-FT is effective with different tuning strategies, showing its robustness.
  • LoRA can outperform full-parameter. When tuning using our BAAI-2k dataset, we find that Shadow-FT (LoRA) can outperform Shadow-FT (full), such as 69.7 vs. 67.4 on Code-3 when tuning Qwen-3-4B. Interestingly, we find that Shadow-FT (LoRA) typically performs better than Shadow-FT (full) on Math-7. However, considering the conventional tuning methods, FT (full) would perform better [40], such as 75.9 vs. 74.8 on Qwen-3-8B. We leave it to future work for further investigation.
@@ 5.1 Ranks in LoRA
We fine-tune the Llama-3.2-1B using LoRA with different ranks (from 4 to 512), and report the average scores after searching learning rates in {5e-5, 1e-4, 2e-4, 5e-4}. As shown in Figure 4, our proposed Shadow-FT (LoRA) can always outperform conventional LoRA with different ranks, demonstrating the robustness. With a larger rank, the conventional LoRA would perform worse, indicating more severe degeneration when tuning the Instruct model [43]. In contrast to that, Shadow-FT (LoRA) can consistently benefit from more parameters (with larger ranks) and achieves better performance. For the results on Llama-3.1-8B, please refer to Appendix B.5.
Figure 3: The average of Math-7, Code-3, and Reasoning-9 for different ranks when tuning Llama-3.2-1B using LoRA. We report the best performance searching learning rates in {5e-5, 1e-4, 2e-4, 5e-4}.

### v3

# Training on Paired Base Model

@@ Abstract
Large language models (LLMs) consistently benefit from further fine-tuning on various tasks. However, we observe that directly tuning the Instruct (i.e., instruction-tuned) models often leads to marginal improvements and even performance degeneration. Notably, paired Base models, the foundation for these Instruct variants, contain highly similar weight values (i.e., less than 2% on average for Llama 3.1 8B). The Base model tends to be a good learner yet a weak backbone without post-training. Therefore, we propose a novel Shadow-FT framework to tune the Instruct models by leveraging corresponding Base models. The key insight is to fine-tune the Base model, and then directly graft the learned weight updates to the Instruct model. Our proposed Shadow-FT introduces no additional parameters, is easy to implement, and significantly improves performance. We conduct extensive experiments on tuning mainstream LLMs, such as Qwen 3 and Llama 3 series, and evaluate them across 19 benchmarks covering coding, reasoning, and mathematical tasks. Experimental results demonstrate that Shadow-FT consistently outperforms conventional full-parameter and parameter-efficient tuning approaches. Further analyses indicate that Shadow-FT can be applied to multimodal large language models (MLLMs) and combined with direct preference optimization (DPO).
@@ 1 Introduction
Large Language Models (LLMs), such as Qwen (Bai et al., 2023), Llama (AI@Meta, 2024), and Gemma (Team et al., 2025a), have demonstrated remarkable performance across diverse disciplines (Zhang et al., 2023; Wang et al., 2024a). Such a strong capability is always attributed to the pre-training on massive data with billions of parameters (Bi et al., 2024; Tao et al., 2024). When applied in real-world scenarios, there are several challenges. The users want the LLMs to follow their instructions helpfully and honestly (Li et al., 2024), which is not covered during the pre-training (Zhang et al., 2023; Liu et al., 2024). Meanwhile, the downstream tasks always involve specific domain knowledge requiring adaptation (Wang et al., 2023; Luo et al., 2024).
@@
To tackle these issues, one predominant approach is further tuning LLMs on desired tasks, including full parameter fine-tuning and parameter-efficient fine-tuning (Liu et al., 2021; Hu et al., 2022). Typically, for each model size, two paired variants are provided: the pretrained base model (denoted as Base) and its instruction-tuned version (denoted as Instruct). The Base model exhibits relatively poor instruction-following ability (i.e., a weak backbone), while the Instruct model performs better. However, we observe that tuning the Instruct models brings marginal improvements and even a performance degeneration. Therefore, how to tune the Instruct model effectively gains increasing importance.
@@
Figure 1: Performance of vanilla SFT (part a), similarity on weights(part b), and the Shadow-FT framework (part c). The progress bars in brown and pink denote the different abilities, the fuller, the better. Based on the SFT dynamics and weight similarity (gap $\sigma$ less than 0.03), we propose to tune the paired Base model and then graft the weight updates onto Instruct model.
@@
In this paper, we first analyze the weights of paired Base and Instruct models considering the relative absolute difference $\sigma$. Fortunately, we find that the weights of Base and Instruct are highly similar. As shown in Figure 1, the gap $\sigma$ is quite low, with an average $\sigma$ of 0.016 for the Llama-3.1-8B model. Intuitively, the contained instruction-following ability of Instruct model disturbs the learning of new knowledge, while Base can avoid it. We further provide a deep analysis to prove this conclusion. Motivated by these, we thus propose a novel Shadow-FT framework to employ the Base model as ’shadow’ of Instruct. The key is to tune the Base for better weight updates and directly graft these updates to Instruct, as they share the same structures.
@@
To evaluate the performance, we conduct extensive experiments tuning mainstream LLMs such as Qwen 3 (Bai et al., 2023) and Llama 3 (AI@Meta, 2024). For the tuning data, we employ the BAAI-Infinity-Instruct Dataset [^1] https://huggingface.co/datasets/BAAI/Infinity-Instruct and extract 2000 samples named as BAAI-2k following (Zhou et al., 2023; Muennighoff et al., 2025). Without the loss of generality, we apply Shadow-FT on full parameter and low-rank settings, and then report the performance on 19 datasets. Experimental results indicate that Shadow-FT consistently outperforms the baselines under various settings, demonstrating its effectiveness and robustness. Further analyses show that Shadow-FT can be applied to MLLMs and combined with DPO for alignment. Our contributions can be concluded as follows:
@@ 2.1 Background
Supervised Fine-tuning (SFT) is a fundamental approach to updating the knowledge of LLMs. Vanilla SFT methods update all the parameters by gradient descent following $W^{+}\leftarrow W+\Delta W$, where $W\in\mathbb{R}^{d_{1}\times d_{2}}$ is an arbitrary weight and $W^{+}$ is the updated variant. To reduce the update costs, LoRA (Hu et al., 2022) introduces a low-rank branch to learn the weight updates following $W^{+}\leftarrow W+AB$, where $A\in\mathbb{R}^{d_{1}\times r},B\in\mathbb{R}^{r\times d_{2}}$ and $r\ll\min\{d_{1},d_{2}\}$. The original weight $W$ is frozen during training, and only the low-rank branch is updated.
@@
Current LLMs typically follow a two-stage training pipeline, including pre-training and post-training. During pre-training, LLMs are trained on massive training data on next token prediction tasks (Brown et al., 2020), and the weights would be released as Base version. The Instruct variant, post-trained upon the Base model, is further aligned with human preference and tuned for reasoning tasks (Ouyang et al., 2022). Therefore, Instruct model performs better than Base model regarding instruction-following ability.
@@ 2.2 Directly Tuning Instruct
However, tuning the Instruct models often leads to marginal improvements and even performance degeneration. Table 1 shows the tuned performance of the Instruct models using the BAAI-2k. We report the average scores of popular benchmarks. Compared to the vanilla Instruct, the tuned version shows marginal improvement, and even degeneration in more cases. Specifically, as shown in Table 1, tuning Qwen-3-4B on the BAAI-2k dataset via conventional LoRA would lead to a drop of 2.6 in Math-7 (from 73.8 to 71.2), 6.8 in Code-3 (from 66.4 to 59.6), and 2.6 in Knowledge-9 (from 63.7 to 61.1). Therefore, how to effectively tune Instruct remains a challenge.
@@ 2.3 Similar Weights: Base & Instruct
where $\sum$ is the element-wise sum and $|\cdot|$ means the absolute operations. The $\sigma$ would be 1 if one is much larger than the other, and be 0 if the two matrices are exactly the same. The smaller the $\sigma$, the more similar the two matrices are. Figure 2 shows the weights of the same layer from Base and Instruct, and also their differences with $\sigma=0.016$. We can find that the gaps are quite small and negligible after zooming in 10x in the z-axis. Please refer to Appendix C for more $\sigma$ regarding various LLMs. In summary, these paired Base and Instruct models are highly similar with $\sigma<0.03$.
@@ 3.1 Shadow-FT
where $Tune$ is the fine-tuning method, such as full-parameter fine-tuning and LoRA. After that, we would like to get the weights updates as the learned knowledge, and directly graft these updates to the Instruct model as:
@@ 3.2 Relation with Task Vectors
Task Vectors aim to represent the ability on tasks as vectors, and are widely used for arithmetic operations on these tasks regarding the same base model (Ilharco et al., 2022). Chat Vector (Huang et al., 2023) extends such an idea to LLMs, which models weight differences between Instruct and Base models as vectors and then adds the vectors to continually pretrained Base models. Specifically, Chat Vector continually pre-trains Llama2 (Touvron et al., 2023) on the Traditional Chinese corpus, and then adds on the chat vectors. Compared to Chat Vector (Huang et al., 2023), the differences are as follows: 1) task: Chat Vector focuses on continual pertaining while Shadow-FT can be applied to board tuning methods, including full-parameter fine-tuning, LoRA, and DPO. 2) motivation: Chat Vector aims to extend the language ability. Shadow-FT aims to tackle the degeneration issue based on the weight similarity.
@@ 4.1 Experimental Setup
For the tuning data, we build BAAI-2k by extracting 2000 samples from BAAI-Infinity-Instruct Dataset  [^2] https://huggingface.co/datasets/BAAI/Infinity-Instruct/tree/main/Gen following (Zhou et al., 2023; Muennighoff et al., 2025). We select the samples with high rewards to ensure the data quality and uniform sampling among all categories for data diversity. Without loss of generality, we tune various LLMs, including Qwen 3 series (Team, 2025) and Llama 3 series (AI@Meta, 2024). Also, we report the results on Gemma-3 series (Team et al., 2025a), Yi series (Young et al., 2024), and Falcon series (Almazrouei et al., 2023) in Section 5.7. We employ LLaMA-Factory (Zheng et al., 2024) for the code base and apply two tuning strategies: full-parameter and LoRA. All experiments are conducted on 8 A100 GPUs. Please refer to Appendix D.1 for detailed hyperparameters.
@@
To evaluate the tuned LLMs on downstream benchmarks, we employ the OpenCompass framework (Contributors, 2023b) and lmdeploy as the acceleration framework (Contributors, 2023a). During inference, we set the cutoff length as 4096 and the batch size as 512. Considering the benchmarks, we select three representative abilities, i.e., mathematical, coding, and commonsense reasoning ability, and report the average scores marked as Math-7, Code-3, and Knowledge-9. Specifically, Math-7 denotes the results of AIME24 MAA (2024), GSM8K (0-shot and 8-shot) Cobbe et al. (2021), MATH Hendrycks et al. (2021b), MATH-500, Minerva_Math Lewkowycz et al. (2022), SVAMP Patel et al. (2021). Code-3 for HumanEval Chen et al. (2021a), HumanEval+ Liu et al. (2023), LiveCodeBench Jain et al. (2024). Knowledge-9 for ARC-challenge Clark et al. (2018), BBH (0-shot and few-shot), DROP Dua et al. (2019), GPQA Diamond Rein et al. (2024), MMLU Hendrycks et al. (2021a), MMLU Pro Wang et al. (2024b), Winogrande ai2 (2019), TheoremQA Chen et al. (2023). To avoid the impact of different prompts, we mainly evaluate under a zero-shot setting. Please refer to Appendix A for more details. For Qwen-3 series, we adapt enable_thinking as false for universal evaluations, and we report pass@k results of both thinking and non-thinking in Appendix D.4.
@@ 4.2 Main Results
Table 1: Performance comparison of different methods tuning popular LLMs. Math-7 denotes the average score of 7 mathematical benchmarks including AIME24, Code-3 for 3 code benchmarks including LiveCodeBench, and Knowledge-9 for 9 commonsense reasoning benchmarks including MMLU Pro. For Math-7 and Code-3, we report the mean value of three runs. We employ the Instruct version and report the final average scores. Please refer to Appendix D.2 for detailed scores.
Model Method Math-7 Code-3 Knowledge-9 Avg.
@@
  • While conventional tuning fails, Shadow-FT performs well in adaptation at the same cost. Across all model sizes and tasks, Shadow-FT consistently outperforms tuning baselines and the vanilla Instruct model. For example, on Qwen-3-4B, Shadow-FT archives an average score of 69.4, which is 3.7 higher than the 65.7 of conventional tuning methods and 1.4 higher than the vanilla Instruct model. The conclusion is consistent on larger models such as Qwen-3-14B. Moreover, Shadow-FT does not introduce any extra training overheads. These consistent gains demonstrate that our proposed Shadow-FT can effectively learn the knowledge contained in training data.
  • Shadow-FT works well under both full-parameter setting and LoRA. For instance, when tuning Qwen-3-4B under full-parameter setting, Shadow-FT achieves 73.7/67.4/64.9 on Math-7/Code-3/Knowledge-9 compared to 72.9/66.4/62.9 of conventional tuning methods. When applying a low-rank setting, Shadow-FT achieves 75.9/69.7/65.0, which is 4.7/10.1/3.9 higher than conventional LoRA. These indicate that Shadow-FT is effective with different tuning strategies, showing its robustness.
  • LoRA can outperform full-parameter. When tuning using our BAAI-2k dataset, we find that Shadow-FT (LoRA) can outperform Shadow-FT (full), such as 69.7 vs. 67.4 on Code-3 when tuning Qwen-3-4B. Interestingly, we find that Shadow-FT (LoRA) typically performs better than Shadow-FT (full) on Math-7. However, considering the conventional tuning methods, FT (full) would perform better (Biderman et al., 2024), such as 75.9 vs. 74.8 on Qwen-3-8B. We leave it to future work for further investigation.
Moreover, we further conduct a case study on Llama-3.1-8B-Instruct. Please refer to Appendix E for more details.
@@ 5.1 Ranks in LoRA
We fine-tune the Llama-3.2-1B using LoRA with different ranks (from 4 to 512), and report the average scores after searching learning rates in {5e-5, 1e-4, 2e-4, 5e-4}. As shown in Figure 4, our proposed Shadow-FT (LoRA) can always outperform conventional LoRA with different ranks, demonstrating the robustness. With a larger rank, the conventional LoRA would perform worse, indicating more severe degeneration when tuning the Instruct model (Yang et al., 2024a). In contrast to that, Shadow-FT (LoRA) can consistently benefit from more parameters (with larger ranks) and achieves better performance. For the results on Llama-3.1-8B, please refer to Appendix D.5.
Figure 3: The average of Math-7, Code-3, and Knowledge-9 for different ranks when tuning Llama-3.2-1B using LoRA. We report the best performance searching learning rates in {5e-5, 1e-4, 2e-4, 5e-4}.